OpenStudy (anonymous):

Ball 2, with a mass of 30kg, is moving to the left at 10m/s. With what velocity should Ball 1, with a mass of 10kg, move to the right and collide with Ball 2, so that Ball 1 rebounds with a velocity of 30m/s, and Ball 2 with a velocity of 10 m/s, after the collision? Assume the collision to be perfectly elastic

OpenStudy (ujjwal):

what is the formula for momentum?

OpenStudy (anonymous):

p=m*v

OpenStudy (ujjwal):

Find total initial momentum of both balls assuming the velocity of ball1 to be v. Find final total momentum of both balls. Linear momentum is always conserved. So, equalize both the total initial and final momentum. you will get v (velocity of ball1)

OpenStudy (anonymous):

okay how do I find the initial momentum do I do 30*10 for ball 2

OpenStudy (anonymous):

and 10*v=30 for ball 1

OpenStudy (anonymous):

you have lost me sorry

OpenStudy (ujjwal):

lost you?

OpenStudy (anonymous):

I don't get what you are saying

OpenStudy (ujjwal):

which part?

OpenStudy (anonymous):

how did you find total initial momentum of each ball?

OpenStudy (ujjwal):

by using the relation you wrote earlier. P=mv

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so then we have 10v+300

OpenStudy (anonymous):

V=30

OpenStudy (ujjwal):

where is the total final momentum? I told you to equalize initial and final momentum. This question involves core knowledge of momentum and you are unaware about the basics. I strongly recommend you to go through the theory part of your text book before you begin solving numerical problems!

OpenStudy (ujjwal):

well the answer turns out to be 30 m/s towards right. It goes like this. Taking left to be +ve direction for velocity. Total initial momentum=30*10+10*v total final momentum=30*(-10)+10*30 equalizing both you get, v=-30m/s That is to say 30 m/s towards right.

OpenStudy (anonymous):

Okay I will go read again thank you for your help.